If 500 kWh cost $57.61, and 1000 kWh cost $108.56, how much does does 713 kWh cost?



Answer by stanbon(75887)

If 500 kWh cost $57.61, and 1000 kWh cost $108.56, how much does does 713 kWh cost?
------------------------
You have two points relating kWh and cost: (500,57.61) and (1000,108.56)
---
Assuming the relation is linear, find the linear equation:
slope = (108.56-57.61)/(1000-500) = 0.1019
---
Intercept: 108.56 = 0.1019*1000 + b
b = 108.56- 101.90
b = 6.66
----------
Equation:
C(x) = 0.1019x + 6.66
----------------------------
So C(713) = 0.1019*713 + 6.66 = $79.31
==========================================
Cheers,
Stan H.